### v1

@@ 4.2 Relation to entanglement wedge cross section in AdS3/CFT2
In this section, we explore a case where the reflected entropy and the entanglement wedge cross section in AdS3/BCFT2 can be viewed as limiting cases of AdS3/CFT2. We analyze the consistency between our work and Jeong:2019xdr by examining this specific case. It appears that the reflected entropy of AdS3/BCFT2 in phase II of the adjacent intervals in Section 3.2 can be seen as a limiting case of the reflected entropy in the connected phase of AdS3/CFT2.
@@
The first condition, which makes $\bar{x}_{2}$ coincide with $\bar{x}_{3}$, reduces one degree of freedom. Motivated by the doubling trick in BCFT (See Figure 6), we place $\bar{x}_{1}$ and $\bar{x}_{4}$ symmetrically with respect to the origin using the second condition, with $\bar{x}_{1}<0$ and $\bar{x}_{4}>0$. In this way, two degrees of freedom remain in total. To address these conditions conveniently, we define new variables as follows:
@@ 5 Conclusion
Our calculations of the Renyi reflected entropy and the entanglement wedge cross section were performed in Euclidean spacetime. An interesting generalization of our work is to introduce Lorentzian time dependence, such as the dynamics of reflected entropy, pseudo entropy, and timelike entanglement entropy Kusuki:2019evw ; Nakata:2020luh ; Doi:2022iyj . In particular, it is important to investigate in detail the analytic continuation of conformal blocks and surface area in the bulk to the Lorentzian signature. Another interesting direction is to include the quantum corrections Faulkner:2013ana ; Engelhardt:2014gca to the entanglement wedge cross section involving backreaction from the cosmic branes. In this situation, quantum extremal surfaces and island-like contributions may come into play Dutta:2019gen ; Chandrasekaran:2020qtn ; Li:2020ceg ; Hayden:2021gno ; Akers:2022max .
@@ Acknowledgements.
We would like to thank Hyun-Sik Jeong for the collaboration at the initial stages of this work. We also thank Jaydeep Kumar Basak and Cynthia Keeler for insightful discussions. This work was supported by the Basic Science Research Program through the National Research Foundation of Korea (NRF) funded by the Ministry of Science, ICT & Future Planning (NRF2021R1A2C1006791) and the Al-based GIST Research Scientist Project grant funded by the GIST in 2024. KYK was also supported by the Creation of the Quantum Information Science R&D Ecosystem (Grant No. 2022M3H3A106307411) through the National Research Foundation of Korea (NRF) funded by the Korean government (Ministry of Science and ICT). B. Ahn was supported by Basic Science Research Program through the National Research Foundation of Korea funded by the Ministry of Education (NRF-2020R1A6A3A01095962, NRF-2022R1I1A1A01064342). S.-E. Bak is supported by the U.S. Department of Energy under grant number DE-SC0019470 and by the Heising-Simons Foundation “Observational Signatures of Quantum Gravity” collaboration grant 2021-2818. M. Nishida was supported by the Basic Science Research Program through the National Research Foundation of Korea (NRF) funded by the Ministry of Education (RS-2023-00245035). B. Ahn, S.-E. Bak and M. Nishida contributed equally to this paper and should be considered as co-first authors.
@@ A.1 Reflected entropy for $m=1$
If the OPE channel takes the form described in Figure 7, the three-point function in the numerator of equation (93) factorizes as follows:
@@
In this phase, there is another contribution from $\langle\sigma_{g_{B}^{-1}g_{A}}(-x_{2})\sigma_{g_{A}^{-1}}(-x_{1})\sigma_{g_{A}}(x_{1})\rangle$ in the chiral CFT, as depicted in Figure 8. As a result, the reflected entropy between two adjacent subregions $A$ and $B$ is given by
@@
with another contribution $\langle\sigma_{g_{B}}(-x_{3})\sigma_{g_{B}g_{A}^{-1}}(x_{2})\sigma_{g_{B}^{-1}}(x_{3})\rangle$ as depicted in Figure 9, then the reflected entropy between two adjacent subregions $A$ and $B$ is obtained as
@@
If the OPE channel doesn’t involve the contribution from the boundary, the reflected entropy in BCFT is the same as the result obtained from the three-point function in CFT (See Figure 10):
@@ A.2 Entanglement wedge cross section for $m=1$
We summarize the entanglement wedge cross section of two intervals $A=[x_{1},x_{2}]$ and $B=[x_{2},x_{3}]$ at a fixed time slice $t=0$ for $m=1$. For $m=1$, the entanglement wedge cross section is obtained by calculating the geodesic length of each case in the geometry without backreaction of the cosmic brane. In principle, this is the same kind of calculation as in Section 3.1. As we discussed in Section 4.1, there are disconnected and connected RT configurations for the subregion $A\cup B$.
@@
For the disconnected RT configuration, there are three possible phases of the entanglement wedge cross section (See Figure 4).
@@
For the connected configuration, $x_{1}$ and $x_{3}$ are two boundary end points of the RT surface. In this case, there is only one phase of the entanglement wedge cross section that extends from the point $x_{2}$ on the boundary and ends at the RT surface. (See Figure 5). The entanglement wedge cross section is given by

### v2

@@ 4.2 Relation to entanglement wedge cross section in AdS3/CFT2
In this section, we explore a case where the reflected entropy and the entanglement wedge cross section in AdS3/BCFT2 can be viewed as limiting cases of AdS3/CFT2. We analyze the consistency between our work and Jeong:2019xdr by examining this specific case. It appears that the reflected entropy of AdS3/BCFT2 in phase II of the adjacent intervals in Sec. 3.2 can be seen as a limiting case of the reflected entropy in the connected phase of AdS3/CFT2.
@@
The first condition, which makes $\bar{x}_{2}$ coincide with $\bar{x}_{3}$, reduces one degree of freedom. Motivated by the doubling trick in BCFT (See Fig. 6), we place $\bar{x}_{1}$ and $\bar{x}_{4}$ symmetrically with respect to the origin using the second condition, with $\bar{x}_{1}<0$ and $\bar{x}_{4}>0$. In this way, two degrees of freedom remain in total. To address these conditions conveniently, we define new variables as follows:
@@ 5 Conclusion
BCFT has been actively studied in the context of entanglement entropy and the black hole information problem. Notably, it plays a central role in reproducing the island prescription within doubly holographic setups and enables modeling black hole evaporation through holographic moving mirror models with time-dependent boundaries Akal:2020twv ; Akal:2021foz ; Akal:2022qei ; BasakKumar:2022stg . By introducing boundaries into CFT, one can explore a broader class of physical situations and observe a richer structure of entanglement, including a wider range of configurations and associated phase transitions. Building on our classification of possible phases and the corresponding $m-1$ corrections, our results can be naturally extended to the moving mirror setup to track how the entanglement wedge cross section evolves as the mirror radiates. In this context, the $m-1$ correction to the reflected entropy may serve as a useful probe of non-trivial phase transitions that occur during holographic black hole evaporation, making this an intriguing direction for future research.
Our calculations of the Renyi reflected entropy and the entanglement wedge cross section were performed in Euclidean spacetime. An interesting generalization of our work is to introduce Lorentzian time dependence, such as the dynamics of reflected entropy, pseudo entropy, and timelike entanglement entropy Kusuki:2019evw ; Nakata:2020luh ; Doi:2022iyj . In particular, it is important to investigate in detail the analytic continuation of conformal blocks and surface area in the bulk to the Lorentzian signature. Another interesting direction is to include quantum corrections Faulkner:2013ana ; Engelhardt:2014gca to the entanglement wedge cross section, taking into account the backreaction from the cosmic branes. In this situation, quantum extremal surfaces and island-like contributions may come into play Dutta:2019gen ; Chandrasekaran:2020qtn ; Li:2020ceg ; Hayden:2021gno ; Akers:2022max .
@@ Acknowledgements.
We would like to thank Hyun-Sik Jeong for the collaboration at the initial stages of this work. We also thank Jaydeep Kumar Basak and Cynthia Keeler for insightful discussions. This work was supported by the Basic Science Research Program through the National Research Foundation of Korea (NRF) funded by the Ministry of Science, ICT & Future Planning (NRF-2021R1A2C1006791) and the Al-based GIST Research Scientist Project grant funded by the GIST in 2024. The work was also supported by the National Research Foundation of Korea (NRF) grant funded by the Korea Government (MSIT) (NRF-2023K2A9A1A01095488). Hospitality at APCTP during the program “Holography 2024: Correlation and Entanglement in Quantum matter” is kindly acknowledged. KYK was also supported by the Creation of the Quantum Information Science R&D Ecosystem (Grant No.2022M3H3A106307411) through the National Research Foundation of Korea (NRF) funded by the Korean government (Ministry of Science and ICT). B. Ahn was supported by Basic Science Research Program through the National Research Foundation of Korea funded by the Ministry of Education (NRF-2020R1A6A3A01095962, NRF-2022R1I1A1A01064342). S.-E. Bak is supported by the U.S. Department of Energy under grant number DE-SC0019470 and by the Heising-Simons Foundation “Observational Signatures of Quantum Gravity” collaboration grant 2021-2818. M. Nishida was supported by the Basic Science Research Program through the National Research Foundation of Korea (NRF) funded by the Ministry of Education (RS-2023-00245035). B. Ahn, S.-E. Bak and M. Nishida contributed equally to this paper and should be considered as co-first authors.
@@ A.1 Reflected entropy for $m=1$
If the OPE channel takes the form described in Fig. 7, the three-point function in the numerator of equation (96) factorizes as follows:
@@
In this phase, there is another contribution from $\langle\sigma_{g_{B}^{-1}g_{A}}(-x_{2})\sigma_{g_{A}^{-1}}(-x_{1})\sigma_{g_{A}}(x_{1})\rangle$ in the chiral CFT, as depicted in Fig. 8. As a result, the reflected entropy between two adjacent subregions $A$ and $B$ is given by
@@
with another contribution $\langle\sigma_{g_{B}}(-x_{3})\sigma_{g_{B}g_{A}^{-1}}(x_{2})\sigma_{g_{B}^{-1}}(x_{3})\rangle$ as depicted in Fig. 9, then the reflected entropy between two adjacent subregions $A$ and $B$ is obtained as
@@
If the OPE channel doesn’t involve the contribution from the boundary, the reflected entropy in BCFT is the same as the result obtained from the three-point function in CFT (See Fig. 10):
@@ A.2 Entanglement wedge cross section for $m=1$
We summarize the entanglement wedge cross section of two intervals $A=[x_{1},x_{2}]$ and $B=[x_{2},x_{3}]$ at a fixed time slice $t=0$ for $m=1$. For $m=1$, the entanglement wedge cross section is obtained by calculating the geodesic length of each case in the geometry without backreaction of the cosmic brane. In principle, this is the same kind of calculation as in Sec. 3.1. As we discussed in Sec. 4.1, there are disconnected and connected RT configurations for the subregion $A\cup B$.
@@
For the disconnected RT configuration, there are three possible phases of the entanglement wedge cross section (See Fig. 4).
@@
For the connected configuration, $x_{1}$ and $x_{3}$ are two boundary end points of the RT surface. In this case, there is only one phase of the entanglement wedge cross section that extends from the point $x_{2}$ on the boundary and ends at the RT surface. (See Fig. 5). The entanglement wedge cross section is given by